Subject: Minutes of User's Committee Meeting

From: Michael Strauss

Submitted: Thu, 14 Dec 1995 09:38:32 -0500

Message number: 31 (previous: 30, next: 32 up: Index)

	Minutes to APO 3.5m User's Committee
	 Phone Conference, December 11, 1995
		Michael Strauss

  Attending: Alan Uomoto, Julie Lutz, Bruce Gillespie, Ed Turner, Lew
Hobbs, Michael Strauss, Rene Walterbos, Chris Stubbs 

The following is an e-mail from Bruce Gillespie describing the current
status of systems at the telescope, with additions based on the
discussions held. 

	Board of Governors annual meeting-
Approved 1996 Operations Budget at $510K, plus funds for replacement
2ndary mirror ($250K) and continuing engineering support. The board
understands that we don't have a hard cost on the 2ndary, and we might
have to come back to them for more (we're planning to get started on
this fast). Deferred hiring of programmer.

 Telescope has generally worked well during the three weeks since the
shutdown.  Pointing is ~4 arcseconds rms, blind tracking OK for few
tens of minutes, and now the offset guider now is working fairly
routinely.  Pointing accuracy seems to be a function of rotation of
secondary.  Seeing and image quality generally near 1 arcsecond (not
that good with DIS).  ChAOS measurements made by Kibblewhite's team
indicate that 20Hz image motion reduced, but still show evidence of
harmonic motion at other frequencies at amplitudes of few tenths
arcseconds--work continuing on tracking down and fixing this.  A
"breathing mode" is seen on ~1 Hertz, due to primary mirror support
hydraulics?  Short exposure image (0.1 sec) taken recently has FWHM of
~.6 arcseconds (this was on a night with good seeing with longer
exposures, around 1"), consistent with Hartmann and profilometry data.
Also 0.7" seeing seen on GRIM evening of December 10.

Stubbs: Are we 100% sure that the secondary is indeed the cause of the
poor seeing? That is, are we sure that we should have it refigured? We
should be so before we commit ourselves? Tests could involve rotating
the secondary 90 degrees, which would take of the order of a week. In
any case, we would not be committed monetarily until summertime.
Strong feeling of the user's committee that one more confirmation is
needed that it is indeed the secondary before committing ourselves. 

Enclosure drive working well since shutdown, cracked wheel problem
under discussion with original manufacturer, L&F Industries; perhaps
we can get them to fix it. We've put the idea into their heads; they
will get back to us. Another ultrasound soon will tell us how long
before these wheels get critical.

		S/W and Remote Observing-
New version of Remark baselined, handles 512^2 guider chip correctly.
Network quality and bandwidth about the same as usual, being
monitored.  Considering extending grace period of data transfer (at
the moment, one has 36 hours to transfer data before it is erased).
We probably have enough disk to get away with this.  One possibility
is that we have a setup like at Keck or KPNO; all data taken are
spooled to tape stored at APO every morning.  Stubbs: Aggressively
investigate what is required for us to make backup copies of all
observations at the site: everyone agreed.
  No progress reported on getting SLIP modem capability in place at
remote stations (not quite true: Princeton has been able to do this!).

DIS is performing nominally, a new calibration spectrum atlas is under
construction and will be placed on the web server when completed.
Andy Silber (UWash) has developed a slit-view CCD (300x400) camera,
which was installed on DIS two weeks ago and tested.  In ~1m
exposures, detects 19th magnitude objects, but has vignetting and
scattered light problems which are being solved, should be operational
in a few weeks as a user facility.

GRIM2 is working well, a reported fix in the exposure time problem
reported recently on the Princeton mail exploder.

Echelle, HRI; no report.

DSC is expected to return to APO next week, having had its dewar (and
CCD) cleaned and leak-tested, plus electronic repairs which are
expected to solve the noise and latent charge problems. No filter
wheel, as far as Bruce knows. 

ChAOS has been used recently in stand-alone mode, some spectacular
results reported.  Laser beacon system being installed at right
Nasmyth, first "beacon" scheduled for next weekend (a great deal of
government approval is needed to shoot lasers into the air).

AMBER camera installation has begun, should continue and be completed
in January.

When shutter for 1024^2 Spectrasource CCD guider failed, replaced
camera with 512^2 camera (which is a backup and also intended for
manual guiding for the Echelle).  Eric Deutsch updated the guide star
finding utility on the web server to work for the smaller format
camera.  512^2 camera had been previously "rehab'd" by the vendor, and
as a guider works very well.  Sensitivity is good enough to close
guide loop on ~17th magnitude objects.  In the richer fields, guide
stars can be found almost at random. This leads to optimism that the
1024^2 camera can be repaired and successfully used (this may be
enough to give a guide star in any field at any PA; if not, we could
put a field lens in front to get another factor of four in field).
Stubbs et al have an engineering run this week to characterize this
well. Camera is being shipped to Spectrasource this week for
evaluation and repair. They may actually change the 1024^2 chip.
Feedback from users is needed on how the guider is working. 

		Calibration Systems-
Have received new Xenon and Krypton arc lamps.  Down-pointing lamps
have been in use for three weeks, exposure times are shorter and now
not necessary to open/close enclosure shutter to take calibrations
(still have to manually open/close mirror covers). Have h/w to now
install three types of arc lamps plus two white light sources
simultaneously, will do this week.  Still have to determine uniformity
of calibrations, and baffling of scattered light.  Plan to add
Calibrations page to web server home page shortly. A real worry is
that the line centers seem to shift depending on which such lamp is on
(a fraction of a pixel is the amplitude of the effect). Flats are also
not as reproducible as we would hope, depending on rotation of the
spectrograph. More efficient baffling would solve this latter problem.
In the meantime, there is a recommendation to take flats over a range
of rotation angles, and average the results. 

		All-sky camera-
  The 10-micron all-sky camera is run routinely by the Observing
Specialists, and Eric Deutsch has written a utility that extracts the
images, does some reduction, and posts these on his web page:
.  A pointer to this has been added to APO's weather web
page--observers can keep tabs on the sky at APO in near real-time.

***********END OF GILLESPIE'S E-MAIL****************

Schedule for next quarter will be out December 15. 

Questions for User's committee from Board, to think about:
  1. Which instrument should JHU build once they become full
  2. Echelle and HRI: Chicago is about to make one further attempt to
get the Echelle going. They are way over the original budget, so if
they fail, they will ask the collaboration to share the costs in this.
Another question: should the HRI be abandoned? Everyone agrees that
the DSC does everything that this would do. Chris Stubbs is thinking
about making a 2048^2 chip into an imager with a big field; this plus
the DSC will cover all the science drivers of the original HRI. The
HRI is on indefinite hold at Chicago anyway (all the effort is going
into the Echelle). User's committee has no objection to dropping the

  An agenda item for next time, from Stubbs: 
   Each institution give a prioritization of the developmental,
enhancement tasks to be done (improving baffling, improving
calibrations, upping observing efficiencies). Further discussion of
how things get fixed once problems are noted.

 Next meeting is January 8. 
APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 31 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to